xx NullExact

NullExact

Bela Szilagyi <bela.szilagyi@jpl.nasa.gov> N T Bishop <N T Bishop <bishont@unisa.ac.za>> Chi Wai Lai <Kevin <laicw@astro.unisa.ac.za>>

August 25 2005

Abstract

1 Introduction

2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Examples 4.6 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References 6 Parameters

11axr0 10Scope: privatex01 REAL

11Description: Constant part of Bondi Radius

11 Range 10 x01 Default: (none)
11-1:1  10amplitude of the perturbation

11

11amplitude  10Scope: privatex01  REAL

11Description: testbed amplitude

11 Range 10 x01 Default: (none)
11*:*  10no restrictions

11

11analyticxfields  10Scope: privatex01  BOOLEAN

11Description: should the analytic values of the fields be computed?

11                         10  x01  Default: no

11

11beta0i  10Scope: privatex01  REAL

11Description: Imaginary part of constant beta0 in linearized solution

11 Range 10 x01 Default: (none)
11*:*  10no restrictions

11

11beta0r  10Scope: privatex01  REAL

11Description: Real part of constant beta0 in linearized solution

11 Range 10 x01 Default: (none)
11*:*  10no restrictions

11

11constantxc1i  10Scope: privatex01  REAL

11Description: Imaginary part of constant C1 in linearized solution

11 Range 10 x01 Default: (none)
11*:*  10no restrictions

11

11constantxc1r  10Scope: privatex01  REAL

11Description: Real part of constant C1 in linearized solution

11 Range 10 x01 Default: (none)
11*:*  10no restrictions

11

11constantxc2i  10Scope: privatex01  REAL

11Description: Imaginary part of constant C2 in linearized solution

11 Range 10 x01 Default: (none)
11*:*  10no restrictions

11

11constantxc2r  10Scope: privatex01  REAL

11Description: Real part of constant C2 in linearized solution

11 Range 10 x01 Default: (none)
11*:*  10no restrictions

11

11errorxmaskxtype  10Scope: privatex01  KEYWORD

11Description: type of mask for error in J

11 Range 10 x01 Default: none
11none  10no mask
11EGxmask  10mask all the active points
11EQxmask  10mask only the equator points
11EVxmask  10mask only the evolution points
11guardxmask  10mask only the guard points

11fxr0  10Scope: privatex01  REAL

11Description: Constant part of Bondi Radius

11 Range 10 x01 Default: 1
110:*  10frequency of the perturbation

11

11omm  10Scope: privatex01  REAL

11Description: frequency

11 Range 10 x01 Default: (none)
11*:*  10no restrictions

11

11testbedxid  10Scope: privatex01  INT

11Description: what testbed should we use?

11 Range 10 x01 Default: (none)
11                         10minkowski, in the natural associated Bondi frame
111  10noise around minkoswski
112  10this is a linear l=2,3 testbed

11

11verbose  10Scope: privatex01  BOOLEAN

11Description: should this thorn be verbose?

11                         10  x01  Default: no

11

11lxinxylm  10Scope: restrictedx01  INT

11Description: l in Ylm

11 Range 10 x01 Default: 2
110:*  10positive

11

11mxinxylm  10Scope: restrictedx01  INT

11Description: m in Ylm

11 Range 10 x01 Default: (none)
11*:*  10must be between -l and +l

11

11nullexactxmaxevolvedarraysize 10Scope: restrictedx01  INT

11Description: The maximum size of evolved grid arrays used by Null2DWave

11 Range 10 x01 Default: 1
111:*  10The size of the generator grid arrays

11

11nullexactxnumarrayevolvedvars 10Scope: restrictedx01  INT

11Description: The maximum number of evolved grid arrays used by NullExact

11 Range 10 x01 Default: 2
112:2  10Is is only the News

11

11linearizedxnews  10Scope: shared from NULLNEWSx01BOOLEAN

7 Interfaces

General Implements: nullexact

Inherits: nullvars nullgrid nullnews nullconstr nullinterp nullevolve

Grid Variables

7.0.1 PRIVATE GROUPS

11  x Group Names x  10x Variable Names xx01 Details x11x
11  realcharfuncsxe  10  x01  compact110
11                         10  wcnxe  x01 description11Error in Evolved Real-valued Characteristic Fields
11                         10  bcnxe  x01 dimensions112
11                         10  x01 distribution11DEFAULT
11                         10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                         10  x01 group type11ARRAY
11                         10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                         10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                         10  x01 timelevels111
11                         10  x01vararrayxsize11NullGrid::Nxradialxpts
11                         10  x01variable type11REAL
11  realconstrxe  10  x01  compact110
11                         10  R01nxe  x01 description11Error in constraints
11                         10  R01sxe  x01 dimensions112
11                         10  x01 distribution11DEFAULT
11                         10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                         10  x01 group type11ARRAY
11                         10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                         10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                         10  x01 timelevels111
11                         10  x01vararrayxsize11NullGrid::Nxradialxpts
11                         10  x01variable type11REAL
11 cmplxcharfuncsxbasicxexsplit 10  x01  compact110
11                         10  jcnr  x01 description11Error in Evolved complex-valued Characteristic Fields (basic set
11                         10  x  x01 description11real and imag part)
11                         10  jcni  x01 dimensions112
11                         10  jcnrxe  x01 distribution11DEFAULT
11                         10  jcnixe  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                         10  x  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                         10  x01 group type11ARRAY
11                         10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                         10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                         10  x01 timelevels111
11                         10  x01vararrayxsize11NullGrid::Nxradialxpts
11                         10  x01variable type11REAL
11cmplxcharfuncsxbasicxexmask10  x01  compact110
11                         10  Mjcnxe  x01 description11Error in Evolved complex-valued Characteristic Fields (basic set)
11                         10  Mucnxe  x01 dimensions112
11                         10  Meth2jcnxe  x01 distribution11DEFAULT
11                         10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                         10  x01 group type11ARRAY
11                         10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                         10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                         10  x01 timelevels111
11                         10  x01vararrayxsize11NullGrid::Nxradialxpts
11                         10  x01variable type11COMPLEX
11 cmplxcharfuncsxbasicxe 10  x01  compact110
11                         10  jcnxe  x01 description11Error in Evolved complex-valued Characteristic Fields (basic set)
11                         10  ucnxe  x01 dimensions112
11                         10  qcnxe  x01 distribution11DEFAULT
11                         10  eth2jcnxe  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                         10  x  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                         10  x01 group type11ARRAY
11                         10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                         10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                         10  x01 timelevels111
11                         10  x01vararrayxsize11NullGrid::Nxradialxpts
11                         10  x01variable type11COMPLEX
11 cmplxcharfuncsxauxxe 10  x01  compact110
11                         10  nucnxe  x01 description11Error in Evolved complex-valued Characteristic Fields (auxiliary set)
11                         10  cbcnxe  x01 dimensions112
11                         10  ckcnxe  x01 distribution11DEFAULT
11                         10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                         10  x01 group type11ARRAY
11                         10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                         10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                         10  x01 timelevels111
11                         10  x01vararrayxsize11NullGrid::Nxradialxpts
11                         10  x01variable type11COMPLEX

11  x Group Names x  10x Variable Names xx01 Details x11x
11 realcharfuncsxexsouth 10  x01  compact110
11                         10  wcsxe  x01 description11Error in Evolved Real-valued Characteristic Fields (south)
11                         10  bcsxe  x01 dimensions112
11                         10  x01 distribution11DEFAULT
11                         10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                         10  x01 group type11ARRAY
11                         10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                         10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                         10  x01 timelevels111
11                         10  x01vararrayxsize11NullGrid::Nxradialxpts
11                         10  x01variable type11REAL
11cmplxcharfuncsxbasicxexsouth10  x01  compact110
11                         10  jcsxe  x01 description11Error in Evolved complex-valued Characteristic Fields (basic set
11                         10  x  x01 description11south)
11                         10  ucsxe  x01 dimensions112
11                         10  qcsxe  x01 distribution11DEFAULT
11                         10  eth2jcsxe  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                         10  x  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                         10  x01 group type11ARRAY
11                         10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                         10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                         10  x01 timelevels111
11                         10  x01vararrayxsize11NullGrid::Nxradialxpts
11                         10  x01variable type11COMPLEX
11 cmplxcharfuncsxauxxexsouth 10  x01  compact110
11                         10  nucsxe  x01 description11Error in Evolved complex-valued Characteristic Fields (auxiliary set
11                         10  x  x01 description11south)
11                         10  cbcsxe  x01 dimensions112
11                         10  ckcsxe  x01 distribution11DEFAULT
11                         10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                         10  x01 group type11ARRAY
11                         10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                         10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                         10  x01 timelevels111
11                         10  x01vararrayxsize11NullGrid::Nxradialxpts
11                         10  x01variable type11COMPLEX
11 newsbr 10 NewsBr x01 compact110
11                         10  x01 description11Real part of Bondi News
11                         10  x01 dimensions112
11                         10  x01 distribution11DEFAULT
11                         10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                         10  x01 group type11ARRAY
11                         10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                         10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                         10  x01 timelevels111
11                         10  x01vararrayxsize112
11                         10  x01variable type11REAL
11 newsbi 10 NewsBi x01 compact110
11                         10  x01 description11Imaginary part of Bondi News
11                         10  x01 dimensions112
11                         10  x01 distribution11DEFAULT
11                         10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                         10  x01 group type11ARRAY
11                         10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                         10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                         10  x01 timelevels111
11                         10  x01vararrayxsize112
11                         10  x01variable type11REAL
11  hmu  10  x01  compact110
11                         10  muh  x01 description11real part of omega - 1
11                         10  EQmuh  x01 dimensions112
11                         10  x01 distribution11DEFAULT
11                         10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                         10  x01 group type11ARRAY
11                         10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                         10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                         10  x01 timelevels111
11                         10  x01vararrayxsize112
11                         10  x01variable type11REAL

11x Group Names x10x Variable Names xx01 Details x11x
11  hjxl  10  x01  compact110
11                 10  Jhxl  x01 description11derivative of J
11                 10  x  x01 description11over x
11                 10  Jhxl  x01 description11evolved values
11                 10  EQJhxl  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize112
11                 10  x01variable type11COMPLEX
11  muxerr  10  x01  compact110
11                 10  muxe  x01 description11error in real part of omega - 1
11                 10  EGmuxe  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize112
11                 10  x01variable type11REAL
11  jxlxerr  10  x01  compact110
11                 10  aJxlxe  x01 description11error in Jxl
11                 10  EGaJxlxe  x01 dimensions112
11                 10  nJxlxe  x01 distribution11DEFAULT
11                 10  EGnJxlxe  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize112
11                 10  x01variable type11COMPLEX
11 newsrxe 10 Newsrxe x01 compact110
11                 10  x01 description11error in real part of News
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize112
11                 10  x01variable type11REAL
11 newsixe 10 Newsixe x01 compact110
11                 10  x01 description11error in imaginary part of News
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize112
11                 10  x01variable type11REAL
11 newsbrxe 10 NewsBrxe x01 compact110
11                 10  x01 description11error in real part of Bondi News
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize112
11                 10  x01variable type11REAL

11x Group Names x10x Variable Names xx01 Details x11x
11 newsbixe 10 NewsBixe x01 compact110
11                 10  x01 description11error in imaginary part of Bondi News
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize112
11                 10  x01variable type11REAL
11 newsbxe 10 NewsBxe x01 compact110
11                 10  x01 description11error in Bondi News
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize112
11                 10  x01variable type11COMPLEX
11 psi4rxe 10 Psi4rxe x01 compact110
11                 10  x01 description11error in real part of Psi4
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize112
11                 10  x01variable type11REAL
11 psi4ixe 10 Psi4ixe x01 compact110
11                 10  x01 description11error in imaginary part of Psi4
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize112
11                 10  x01variable type11REAL
11 psi4xe 10 Psi4xe x01 compact110
11                 10  x01 description11error in Psi4
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize112
11                 10  x01variable type11COMPLEX
11 ubondixe 10 uBondixe x01 compact110
11                 10  x01 description11error in Bondi time associated with the Bondi News
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize112
11                 10  x01variable type11REAL

11x Group Names x10x Variable Names xx01 Details x11x
11 anaxjxl 10 anaxJxl x01 compact110
11                 10  x01 description11analytic Jxl
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize112
11                 10  x01variable type11COMPLEX
11 numxjxl 10 numxJxl x01 compact110
11                 10  x01 description11derivative of J
11 10 x x01 description11over x
11                 10  x01 description11numeric values
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize112
11                 10  x01variable type11COMPLEX
11 anaxmu 10 anaxmu x01 compact110
11                 10  x01 description11analytic Omega
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize112
11                 10  x01variable type11REAL
11 anaxnews 10 anaxNews x01 compact110
11                 10  x01 description11analytic complex-valued News (basic set)
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize112
11                 10  x01variable type11COMPLEX
11 anaxpsi4 10 anaxPsi4 x01 compact110
11                 10  x01 description11analytic complex-valued Psi4 (basic set)
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize112
11                 10  x01variable type11COMPLEX
11  anaxj  10  x01  compact110
11                 10  anaxjcn  x01 description11Error in Evolved complex-valued Characteristic Fields J
11                 10  anaxjcs  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize11NullGrid::Nxradialxpts
11                 10  x01variable type11COMPLEX

11x Group Names x10x Variable Names xx01 Details x11x
11  anaxu  10  x01  compact110
11                 10  anaxucn  x01 description11Error in Evolved complex-valued Characteristic Fields U
11                 10  anaxucs  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize11NullGrid::Nxradialxpts
11                 10  x01variable type11COMPLEX
11  anaxq  10  x01  compact110
11                 10  anaxqcn  x01 description11Error in Evolved complex-valued Characteristic Fields Q
11                 10  anaxqcs  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize11NullGrid::Nxradialxpts
11                 10  x01variable type11COMPLEX
11 dotnewsxmol 10 dotNewsxMoL x01  compact110
11                 10  x01 description11real dotNews for MoL - this is the real part of Psi4
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels112
11                 10  x01vararrayxsize112
11                 10  x01variable type11REAL
11 newsxmol 10 NewsxMoL x01 compact110
11                 10  x01 description11real News for MoL - this is the real part of integral Psi4
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels112
11                 10  x01vararrayxsize112
11                 10  x01variable type11REAL
11 errnewsxmol 10 errNewsxMoL x01  compact110
11                 10  x01 description11difference between News from MoL integration of Psi4 and evolved News
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize112
11                 10  x01variable type11REAL
11  ylmxgroup  10  x01  compact110
11                 10  Ylmx2  x01 description11Ylmxs by patch
11                 10  Ylmx1  x01 dimensions112
11                 10  Ylmx0  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize112
11                 10  x01variable type11COMPLEX

11x Group Names x10x Variable Names xx01 Details x11x
11 wtexactxmask 10 x01 compact110
11                 10 WTexactxmaskn x01 description11Worldtube mask
11                 10 WTexactxmasks x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01variable type11INT

11

8 Schedule This section lists all the variables which are assigned storage by thorn PITTNullCode/NullExact. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage x

11Always:  11Conditional:
11NewsxMoL[2] dotNewsxMoL[2] errNewsxMoL11Ylmxgroup WTexactxmask
11x  11NullSHRExtract::xxwt[2] NullSHRExtract::WTxr0
11x  11NullSHRExtract::jxwt[2] NullSHRExtract::betaxwt[2] NullSHRExtract::wxwt[2]
11x  11NullSHRExtract::qxwt[2] NullSHRExtract::uxwt[2] NullSHRExtract::uxxxwt
11x  11NullSHRExtract::jxl NullSHRExtract::betaxl NullSHRExtract::uxl NullSHRExtract::wxl

11x  11x

11

Scheduled Functions

CCTKxINITIAL (conditional) x nullexactxprecalcylm x precalculate spherical harmonics

11x11Before: 11nullevolxboundaryinit
11 x 11x  11nullevolxinitial
11 x 11Language:11fortran
11 x 11Type: 11function

11

CCTKxANALYSIS (conditional) x nullexactxerrorxcmplx x error in characteristic evolution variables x

11x11Language:11fortran
11 x 11Storage: 11cmplxcharfuncsxbasicxexsplit
11 x 11x  11cmplxcharfuncsxbasicxexmask
11 x 11x  11cmplxcharfuncsxbasicxe
11 x 11x  11cmplxcharfuncsxbasicxexsouth
11 x 11Triggers: 11cmplxcharfuncsxbasicxexsplit
11 x 11x  11cmplxcharfuncsxbasicxexmask
11 x 11x  11cmplxcharfuncsxbasicxe
11 x 11x  11cmplxcharfuncsxbasicxexsouth
11 x 11Type: 11function

11

NullEvolxInitial (conditional) x nullexactxinitial x exact initial data for j x

11x11After: 11nullevolxinitialdata
11 x 11Before: 11nullevolxinitialslice
11 x 11Language:11fortran
11 x 11Type: 11function

11

NullEvolxInitial x nullexactxinitmol x initial data x

11x11After: 11nullevolxinitialslice
11 x 11Language:11fortran
11 x 11Options: 11global
11 x 11Type: 11function

11

MoLxRegister x nullexactxmolregister x register variables for mol x

11x11Language:11c
11 x 11Type: 11function

11

MoLxCalcRHS x nullexactxdotnewsmol x calculate the integral of psi4

11x11Language:11fortran
11 x 11Sync:  11dotnewsxmol
11 x 11Type: 11function

11

CCTKxANALYSIS (conditional) x nullexactxerrorxnewsb x error in characteristic evolution variables x

11x11Language:11fortran
11 x 11Storage: 11newsbxe
11 x 11x  11ubondixe
11 x 11x  11newsbixe
11 x 11x  11newsbrxe
11 x 11x  11newsbr
11 x 11x  11newsbi
11 x 11x  11newsrxe
11 x 11x  11newsixe
11 x 11x  11psi4rxe
11 x 11x  11psi4ixe
11 x 11x  11psi4xe
11 x 11x  11muxerr

11 x 11x  11hmu

11 x 11x  11jxlxerr

11 x 11x  11hjxl

11 x 11Triggers: 11newsbxe
11 x 11x  11ubondixe
11 x 11x  11newsbixe
11 x 11x  11newsbrxe
11 x 11x  11newsbr
11 x 11x  11newsbi
11 x 11x  11newsrxe
11 x 11x  11newsixe
11 x 11x  11psi4rxe
11 x 11x  11psi4ixe
11 x 11x  11psi4xe
11 x 11x  11muxerr

11 x 11x  11hmu

11 x 11x  11jxlxerr

11 x 11x  11hjxl

11 x 11Type: 11function

11

CCTKxANALYSIS (conditional) x nullexactxerrorxreal x error in characteristic evolution variables x

11x11Language:11fortran
11 x 11Storage: 11realcharfuncsxe
11 x 11x  11realcharfuncsxexsouth
11 x 11Triggers: 11realcharfuncsxe
11 x 11x  11realcharfuncsxexsouth
11 x 11Type: 11function

11

CCTKxANALYSIS (conditional) x nullexactxerrorxconstr x error in constraints

11x11After: 11nullconstrxdriver
11 x 11Language:11fortran
11 x 11Storage: 11realconstrxe
11 x 11Triggers: 11realconstrxe
11 x 11Type: 11function

11

CCTKxANALYSIS (conditional) x nullexactxerrorxcmplxaux x error in characteristic evolution variables x

11x11Language:11fortran
11 x 11Storage: 11cmplxcharfuncsxauxxe
11 x 11x  11cmplxcharfuncsxauxxexsouth
11 x 11Triggers: 11cmplxcharfuncsxauxxe
11 x 11x  11cmplxcharfuncsxauxxexsouth
11 x 11Type: 11function

11

CCTKxANALYSIS (conditional) x nullexactxanafields x error in characteristic evolution variables x

11x11Language:11fortran
11 x 11Storage: 11anaxnews
11 x 11x  11anaxmu
11 x 11x  11anaxpsi4
11 x 11x  11anaxj
11 x 11x  11anaxq
11 x 11x  11anaxu
11 x 11x  11anaxjxl
11 x 11x  11numxjxl
11 x 11Triggers: 11anaxnews
11 x 11x  11anaxmu
11 x 11x  11anaxpsi4
11 x 11x  11anaxj
11 x 11x  11anaxq
11 x 11x  11anaxu
11 x 11x  11anaxjxl
11 x 11x  11numxjxl
11 x 11Type: 11function

11

NullEvolxBoundaryInit (conditional) x nullexactxboundary x exact boundary data for the characteristic evolution variables x

11x11Language:11fortran
11 x 11Type: 11function

11

NullEvolxBoundaryInit (conditional) x nullexactxboundarypast x exact boundary data for the characteristic evolution variables, past level x

11x11Language:11fortran
11 x 11Type: 11function

NullEvolxBoundary (conditional) x nullexactxboundary x exact boundary data for the characteristic evolution variables x

11x11Language:11fortran
11 x 11Type: 11function